an object is projected with speed 20m/s at an angle 60 with vertical. how much time it will take to cross 2 equalpoles of height 4m
 

Grade:11

1 Answers

Khimraj
5 years ago
time will equal to 2 times of time taken from max height to height of 4m.
max height = u2sin260/2g = 400*3/(4*20) = 15m
then H-4 = 15-4 = 11
then t = 2*(\sqrt{11*2/10}) = 2\sqrt{11/5}s
Hope it clears.